# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v1

@@ Abstract
The Probably Approximately Correct (PAC) Privacy framework [1] provides a powerful instance-based methodology for certifying privacy in complex data-driven systems. However, existing PAC Privacy algorithms rely on a Gaussian mutual information upper bound. We show that this is in general too conservative: the upper bound obtained by these algorithms is tight if and only if the perturbed mechanism output is jointly Gaussian with independent Gaussian noise. To address the inefficiency inherent in the Gaussian-based approach, we introduce Residual-PAC (RPAC) Privacy, an $f$-divergence-based measure that quantifies the privacy remaining after adversarial inference. When instantiated with Kullback–Leibler divergence, Residual-PAC Privacy is governed by conditional entropy. Moreover, we propose Stackelberg Residual-PAC (SR-PAC) privatization mechanisms for RPAC Privacy, a game-theoretic framework that selects optimal noise distributions through convex bilevel optimization. Our approach achieves tight privacy budget utilization for arbitrary data distributions. Moreover, it naturally composes under repeated mechanisms and provides provable privacy guarantees with higher statistical efficiency. Numerical experiments demonstrate that SR-PAC certifies the target privacy budget while consistently improving utility compared to existing methods.
@@ 1 Introduction
Artificial intelligence systems now operate across increasingly critical domains—from healthcare diagnostics and autonomous transportation to personalized financial services and national security infrastructure. While demonstrating powerful capabilities, these data-driven systems introduce significant privacy risks as their training and operational data becomes increasingly sensitive and complex. Individual information can be inadvertently exposed through seemingly harmless outputs. Given the expanding scale of modern data pipelines, robust and reliable privacy guarantees have become essential rather than merely desirable.
The past two decades have witnessed the emergence of numerous privacy definitions and frameworks designed to address information leakage risks. Among these, Differential Privacy (DP) [2] stands as perhaps the most influential, providing strong worst-case guarantees against adversarial inference. In its canonical formulation, DP quantifies the maximum change in output probabilities induced by modifying a single data point, thereby ensuring indistinguishability between neighboring datasets. This theoretical foundation has inspired substantial research and practical deployment in tools used by major organizations (e.g., Apple [3]) and governments (e.g., US Census Bureau [4]). Other privacy notions, such as Maximal Leakage [5, 6], mutual information-based criteria, such as Mutual Information Differential Privacy (MI-DP) [7], and Fisher information-based metric [8, 9, 10], have been proposed to capture different adversarial models and to offer alternative trade-offs between privacy and utility.
@@
Provable privacy guarantees for modern data‐processing algorithms face two major obstacles. First, worst‐case frameworks like DP require computing global sensitivity, which is generally intractable (NP-hard [11]). In addition, computing the optimal privacy bound of DP under composition is, in general, a #P-complete under composition task [12]. Likewise, enforcing MI-DP requires a white-box analysis of every possible input distribution, which is seldom available. In practice, finding the minimal noise needed to meet a target guarantee is intractable for most real-world algorithms, especially when the effect of each operation on privacy is unclear. Empirical or simulation-based methods (e.g., testing resistance to membership inference [13]) address specific threats but lack rigorous, adversary-agnostic assurances. Bridging this gap requires a new, broadly applicable framework that can quantify and enforce privacy risk without relying on ad-hoc sensitivity analyses.
A promising alternative has recently emerged: the Probably Approximately Correct (PAC) Privacy framework [1]. Drawing inspiration from PAC learning theory [14], PAC Privacy fundamentally redefines the privacy objective as the information-theoretic hardness of reconstructing sensitive data given arbitrary information disclosure processing. PAC Privacy shifts the paradigm from indistinguishability to inference impossibility by characterizing the probability that any adversary—regardless of strategy or computational power—can accurately infer private data under a specified reconstruction criterion.
At the heart of PAC Privacy lies the concept of PAC Advantage Privacy (PAC-AP), which extends the framework to arbitrary adversarial inference tasks by quantifying the adversary’s improvement in success probability through an $f$-divergence measure. When the $f$-divergence is instantiated as the Kullback–Leibler (KL) divergence, PAC-AP specializes to a mutual information-based criterion, providing a natural and information-theoretic characterization of privacy risk. A closed-form upper bound on mutual information—derived from the maximum entropy property of the Gaussian distribution—serves as the foundation for automatic PAC privatization algorithms, which operate via end-to-end black-box simulations. This approach circumvents the need for direct mutual information computation and enables practitioners to certify privacy risk with high statistical confidence, even when the internal structure of the data processing mechanism is unknown or intractable. In addition, PAC Privacy also enjoys elegant composition properties: when the composed mechanisms are independent, the total privacy loss can be bounded by the sum of the individual mutual information budgets.
However, the upper bound provided by automatic PAC Privacy algorithms (see Theorem 3 and Corollary 2 of [1], and Theorem 1 of [15]) is generally conservative. In particular, for a given privacy budget, the mutual information achieved under the noise distribution constructed by these algorithms is strictly less than the designated privacy budget unless the mechanism output is Gaussian and the noise is independent, zero-mean Gaussian. The privacy budget is exactly attained—i.e., the true mutual information matches the privacy bound—if and only if these Gaussian conditions hold. As a result, this conservativeness leads to a “waste” of privacy budget.
To maximize privacy budget efficiency, PAC Privacy defines optimal perturbation (Definition 9 of [1]) as the noise distribution that minimizes utility loss (e.g., $\ell_{2}$ norm or noise power) while ensuring the mutual information remains within the privacy budget. Crucially, this mutual information constraint intricately links the noise distribution to both the data distribution and mechanism, often making the identification of optimal noise intractable, especially when the joint distribution lacks a closed-form or exhibits complex dependencies.
We address the limitations imposed by the conservativeness of automatic PAC Privacy algorithms. First, we introduce the notion of Residual-PAC Privacy, using $f$-divergence, to quantify the privacy that remains after information has been leaked by a data processing mechanism. While PAC-AP measures the amount of privacy loss incurred by a given mechanism, our Residual-PAC Privacy framework provides a complementary perspective by characterizing the remaining privacy guarantee that persists despite adversarial inference. When $f$-divergence is instantiated as KL divergence, our Residual-PAC Privacy is fully characterized by the conditional entropy up to a known constant that does not depend on the mechanism or the applied noise.
We formally characterize the conservativeness from the automatic PAC Privacy algorithms. The closed-form upper bound obtained by Theorem 3 of [1] forms the theoretical foundation of the automatic PAC Privacy algorithms. Algorithm 1 of [1] leverages this result to perform end-to-end automatic privatization, systematically constructing a Gaussian noise distribution that implements any designated privacy budget as an upper bound on mutual information. Importantly, we show that the Gaussian noise distribution constructed by Algorithm 1 of [1] is the unique solution to a constrained optimization problem that seeks to minimize the noise magnitude (specifically, the trace of the covariance matrix of the Gaussian noise) subject to the constraint that the resulting Gaussian mutual information does not exceed the prescribed privacy budget. Thus, the PAC private mechanism derived by Algorithm 1 of [1] not only enforces the desired privacy bound, but does so in an optimally noise-efficient manner under the Gaussian assumption.
@@
Inspired by this observation, we propose a novel approach, termed Stackelberg Residual-PAC (SR-PAC), for automating Residual-PAC privatization by formulating the problem as a Stackelberg game, without requiring any white-box characterization of the conditional entropy. In this framework, the leader selects a noise distribution to perturb the mechanism by minimizing the magnitude of the noise or perturbation. The follower then chooses a stochastic inference strategy to recover the sensitive data, seeking to minimize the expected log score function of the strategy, with the constraint that the expected log score equals the privacy budget. The privacy budget in Residual-PAC privacy directly translates to a mutual information bound via the simple relation: mutual information equals data entropy minus the privacy budget. When the entire probability space is considered, this bilevel optimization problem reduces to a convex program. We rigorously prove that the mixed-strategy Stackelberg equilibrium of this game yields the optimal noise distribution, ensuring that the conditional entropy of the perturbed mechanism precisely attains the specified privacy budget.
We demonstrate that, in general, the SR-PAC approach achieves more efficient privacy budget utilization than the automatic PAC Privacy algorithms, resulting in less conservative privacy guarantees for the same privacy budget, except in the special case where the overall distribution of the perturbed mechanism output is Gaussian, in which case both methods coincide and tightly implement the privacy budget. Besides inheriting PAC Privacy’s additive composition property, SR PAC attains tighter privacy bounds by solving a convex optimization problem. Moreover, when the utility loss function is convex, our method admits an explicit convex optimization formulation, enabling the computation of optimal perturbations, which is a challenge that remained open in the original PAC Privacy framework. Finally, we validate these theoretical advances through numerical experiments, which consistently confirm the superior performance and practical advantages of SR-PAC over PAC Privacy.
@@ 1.1 Related Work
Privacy Quantification Notions. Quantitative notions of privacy leakage have been extensively studied across a variety of contexts, leading to mathematically rigorous frameworks for assessing the amount of sensitive information that can be inferred by adversaries. Differential privacy (DP) and its variants have become the gold standard for formal privacy guarantees, with the original definitions by Dwork et al. [16, 2] formalizing privacy loss through bounds on the distinguishability of outputs under neighboring datasets. Variants such as concentrated differential privacy (CDP) [17, 18], zero-concentrated DP (zCDP) [19], and R’enyi differential privacy (RDP) [20] have further extended this framework by parameterizing privacy loss with different statistical divergences (e.g., R’enyi divergence), thereby enhancing flexibility in privacy accounting, especially for compositions and adaptive mechanisms. Information-theoretic measures provide alternative and complementary approaches for quantifying privacy loss. For instance, mutual information has been used to analyze privacy leakage in a variety of settings [21, 7], with $f$-divergence and Fisher information offering finer-grained or context-specific metrics [1, 8, 9, 10]. These frameworks help to bridge the gap between statistical risk and adversarial inference, and are closely connected to privacy-utility trade-offs in mechanism design. Maximal leakage, hypothesis testing privacy, and other relaxations further broaden the analytic toolkit for measuring privacy risk.
Privacy-Utility Trade-off Balancing the trade-off between privacy and utility is a central challenge in the design of privacy-preserving mechanisms. This challenge is frequently formulated as an optimization problem [22, 23, 24, 25, 26, 27, 28, 29, 30]. For example, Ghosh et al. [25] demonstrated that the geometric mechanism is universally optimal for differential privacy under certain loss-minimizing criteria in Bayesian settings, while Lebanon et al. [22] and Alghamdi et al. [29] studied utility-constrained optimization. Gupte et al. [26] modeled the privacy-utility trade-off as a zero-sum game between privacy mechanism designers and adversaries, illustrating the interplay between optimal privacy protection and worst-case loss minimization.
Simulation-Based and Mutual Information Approaches Recent work has shifted toward instance-based and data-driven privacy frameworks that empirically measure information leakage for black-box mechanisms. PAC Privacy [1] and its improvements aim to automatically certify privacy loss by empirically estimating $f$-divergence and mutual information through repeated simulation and sampling. These approaches are especially valuable for complex mechanisms such as deep neural networks, where closed-form analysis or classical sensitivity calculations are infeasible or overly conservative. However, the mutual information constraint in PAC Privacy introduces substantial analytic and computational complexity, as it depends jointly on the data distribution, the mechanism, and the candidate noise distribution. In practice, the framework relies on propose-and-verify procedures, which may be conservative in the absence of tractable optimization.
@@
2 Preliminaries: the PAC Privacy Framework
PAC privacy framework [1] considers the following general privacy problem. A sensitive input $X$ is drawn from a distribution $\mathcal{D}$, which may be unknown or inaccessible. Each private data point $x$ is defined over some measurable domain $\mathcal{X}^{\dagger}$, and the dataset $X\in\mathcal{X}$. There is a data processing (possibly randomized) mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}\subset\mathbb{R}^{d}$, where $\mathcal{Y}$ is measurable. The central privacy concern is determining whether an adversary can accurately estimate the true input via an estimate $\widetilde{X}$ based on the observation $Y=\mathcal{M}(X)$, meeting some predefined success criterion captured by an indicator function $\rho$. PAC Privacy is formally defined as follows.
@@ 2 Preliminaries: the PAC Privacy Framework
For a data processing mechanism $\mathcal{M}$, given some data distribution $\mathsf{D}$, and a measure function $\rho(\cdot,\cdot)$, we say $\mathcal{M}$ satisfies $(\delta,\rho,\mathsf{D})$-PAC Privacy if the following experiment is impossible:
A user generates data $X$ from distribution $\mathsf{D}$ and sends $\mathcal{M}(X)$ to an adversary. The adversary who knows $\mathsf{D}$ and $\mathcal{M}$ is asked to return an estimation $\widetilde{X}\in\mathcal{X}$ on $X$ such that with probability at least $(1-\delta)$, $\rho(\widetilde{X},X)=1$.
@@
A mechanism $\mathcal{M}$ is termed $(\Delta_{f}\delta,\rho,\mathsf{D})$ PAC-advantage private if it is $(\delta,\rho,\mathsf{D})$ PAC private and
$\Delta_{f}\delta\equiv\mathcal{D}_{f}(\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}})=\delta^{\rho}_{o}f(\frac{\delta}{\delta^{\rho}_{o}})+(1-\delta^{\rho}_{o})f(\frac{1-\delta}{1-\delta^{\rho}_{o}}).$
@@
Here, PAC Advantage Privacy is defined on top of PAC Privacy and quantifies the amount of privacy loss incurred from releasing $\mathcal{M}(X)$, measured as the prior-expected posterior advantage using $f$-divergence.
@@ 2.1 Automatic PAC Privatization Algorithms
PAC Privacy enables the automatic privatization for arbitrary black box mechanisms by bounding the mutual information between private data and the released output.
@@
$\mathtt{MI}(x;w)\equiv\mathcal{H}(x)-\mathcal{H}(x|w)=\mathcal{D}_{KL}(\mathsf{P}_{x,w}\|\mathsf{P}_{x}\otimes\mathsf{P}_{w}),$
@@
This approach supports simulation based privatization without requiring the worst-case adversarial analysis, such as sensitivity computation. In this section, we present the main theorems and algorithms underlying automatic PAC privatization as introduced in [1] (hereafter “Auto-PAC”) and the efficiency-improved version proposed in [15] (hereafter “Efficient-PAC”).
We start by introducing Auto-PAC. Consider a deterministic data processing mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where the output norm is uniformly bounded: $\|\mathcal{M}(X)\|_{2}\leq r$ for all $X$. To ensure PAC Privacy, the mechanism is perturbed by Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$, and the privacy leakage is quantified via mutual information between the input $X$ and the noisy output $\mathcal{M}(X)+B$. For any deterministic mechanism $\mathcal{M}$ and noise $B$ according to any distribution, define
@@
where $\Sigma_{\mathcal{M}(X)}$ and $\Sigma_{B}$ are the covariances of $\mathcal{M}(X)$ and $B$.
@@
For an arbitrary deterministic mechanism $\mathcal{M}$ and Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$, the mutual information $\mathtt{MI}(\cdot;\cdot)$ satisfies
@@
Algorithm 1 Auto-PAC [1]
0:  deterministic mechanism $M$, dataset $\mathcal{D}$, sample size $m$, variance floor $c$, mutual information quantities $\beta^{\prime}$ and $v$.

### v2

@@ Abstract
The Probably Approximately Correct (PAC) Privacy framework [46] provides a powerful instance-based methodology to preserve privacy in complex data-driven systems. Existing PAC Privacy algorithms (we call them Auto-PAC) rely on a Gaussian mutual information upper bound. However, we show that the upper bound obtained by these algorithms is tight if and only if the perturbed mechanism output is jointly Gaussian with independent Gaussian noise. We propose two approaches for addressing this issue. First, we introduce two tractable post-processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, the result still leaves "wasted" privacy budget. To address this issue more fundamentally, we introduce Residual-PAC (R-PAC) Privacy, an $f$-divergence-based measure to quantify privacy that remains after adversarial inference. To implement R-PAC Privacy in practice, we propose a Stackelberg Residual-PAC (SR-PAC) privatization mechanism, a game-theoretic framework that selects optimal noise distributions through convex bilevel optimization. Our approach achieves efficient privacy budget utilization for arbitrary data distributions and naturally composes when multiple mechanisms access the dataset. Through extensive experiments, we demonstrate that SR-PAC obtains consistently a better privacy-utility tradeoff than both PAC and differential privacy baselines.
@@ 1 Introduction
Machine-learning models power critical applications—from medical diagnosis to autonomous vehicles—yet their outputs can inadvertently expose sensitive training data. As pipelines grow in scale and complexity, practitioners need rigorous, scalable privacy guarantees that go beyond ad-hoc testing. Over the past two decades, formal privacy frameworks have proliferated. Differential Privacy (DP) [13] (and its variants such as Rényi DP [33]) delivers input-independent worst-case indistinguishability by bounding output shifts from single-record changes. Alternative information-theoretic definitions, such as mutual-information DP [10], Fisher-information bounds [16, 23, 21], and Maximal Leakage [26, 39], provide complementary guarantees and offer alternative trade-offs between privacy and utility.
@@
Nevertheless, provable privacy guarantees for modern data-processing algorithms remains a challenge. First, worst-case frameworks like DP require computing global sensitivity, which is generally NP-hard [47]. Moreover, computing the optimal privacy bound of DP under composition is, in general, a #P-complete [34]. In practice, finding the minimal noise needed to meet a target guarantee is intractable for most real-world algorithms, especially when the effect of each operation on privacy is unclear. On the other hand, empirical or simulation-based methods (e.g., testing resistance to membership inference [42]) address specific threats but lack rigorous, adversary-agnostic assurance. Bridging this gap requires a new, broadly applicable framework that can quantify and enforce privacy risk without relying on sensitivity.
A promising alternative has recently emerged: the Probably Approximately Correct (PAC) Privacy framework [46]. PAC Privacy shifts from indistinguishability-based guarantees to an operational notion that measures the information-theoretic hardness of reconstructing sensitive data. It is defined by an impossibility-of-inference guarantee for a chosen adversarial task and data prior, and the framework provides algorithms that enforce tractable mutual-information upper bounds to certify this guarantee. This approach enables automatic privatization via black-box simulation, and enjoys additive composition bounds and automatic privacy budget implementations for adaptive sequential compositions of mechanisms with arbitrary interdependencies. Notably, PAC Privacy often requires only $O(1)$ noise magnitude to achieve its privacy guarantees—independent of the output dimension—whereas differential privacy’s worst-case, input-independent noise magnitude scales as $\Theta(\sqrt{d})$ for a $d$-dimensional release.
However, existing PAC privacy algorithms are fundamentally conservative. In particular, we show (Proposition 1) that Auto-PAC achieves the designated privacy budget exactly if and only if the perturbed mechanism output is jointly Gaussian with independent Gaussian noise, a restrictive condition rarely met in practice. Consequently, Auto-PAC will in general make inefficient use of the privacy budget.
We address this limitation of Auto-PAC in two ways. First, working within the general PAC Privacy framework, we develop two tractable post-processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, even these methods fail to fully close the privacy budget gap. To address this issue more fundamentally, we introduce the notion of Residual-PAC Privacy (R-PAC privacy). Unlike PAC privacy, which aims to directly bound mutual information, R-PAC privacy focuses instead on quantifying privacy budget remaining after information has been leaked by a data processing mechanism, using $f$-divergence to this end. When $f$-divergence is instantiated as Kullback–Leibler (KL) divergence, we show that Residual-PAC Privacy is fully characterized by the conditional entropy up to a known constant that does not depend on the mechanism or the applied noise.
@@
As a practical instantiation of R-PAC, we propose a novel Stackelberg Residual-PAC (SR-PAC) framework. SR-PAC formulates the problem of adding noise given a privacy budget as a Stackelberg game in which the leader selects a noise distribution with the goal of minimizing the magnitude of the perturbation, while the follower chooses a stochastic inference strategy to recover the sensitive data. We show that when the entire probability space is considered, the resulting bilevel optimization problem becomes a convex program. Moreover, we prove that the mixed-strategy Stackelberg equilibrium of this game yields the optimal noise distribution, ensuring that the conditional entropy of the perturbed mechanism precisely attains the specified privacy budget. Finally, we use extensive experimental evaluation to demonstrate that the proposed SR-PAC privacy framework consistently outperforms both PAC-privacy and differential privacy baselines.
In summary, our main contributions are as follows:
  • We characterize the conservativeness of Auto-PAC [46, 43], showing that it arises from the gap between the Gaussian surrogate bound and the true non-Gaussian mutual information of the privatized mechanism.
  • We propose two computationally tractable approaches to reduce this gap: one based on the Donsker-Varadham representation (Theorem 3) and another based on the sliced Wasserstein distances (Theorem 4), both providing sample-efficient non-Gaussianity corrections.
  • We propose a novel privacy framework, Residual-PAC (R-PAC), to quantify the portion of privacy that remains rather than the amount leaked. This offers a complementary perspective to PAC privacy, and enables efficient computation of tight privacy bounds.
  • We present an automatic privatization algorithm, Stackelberg R-PAC (SR-PAC), to efficiently compute noise distributions for a given privacy budget. SR-PAC algorithm achieves tight budget utilization, can operate with only black-box access via Monte Carlo simulation, and adaptively concentrates noise in privacy-sensitive directions while preserving task-relevant information.
@@ 1.1 Related Work
Privacy Quantification Notions. Quantitative notions of privacy leakage have been extensively studied across a variety of contexts, leading to mathematically rigorous frameworks for assessing the amount of sensitive information that can be inferred by adversaries. Differential privacy (DP) and its variants have become the gold standard for formal privacy guarantees, with the original definitions by Dwork et al. [14, 13] formalizing privacy loss through bounds on the distinguishability of outputs under neighboring datasets. Variants such as concentrated differential privacy (CDP) [5, 15], zero-concentrated DP (zCDP) [4], and Rényi differential privacy (RDP) [33] have further extended this framework by parameterizing privacy loss with different statistical divergences (e.g., R’enyi divergence), thereby enhancing flexibility in privacy accounting, especially for compositions and adaptive mechanisms. Information-theoretic measures provide alternative and complementary approaches for quantifying privacy loss. For instance, mutual information has been used to analyze privacy leakage in a variety of settings [7, 10], with $f$-divergence and Fisher information offering finer-grained or context-specific metrics [46, 16, 23, 21]. These frameworks help to bridge the gap between statistical risk and adversarial inference, and are closely connected to privacy-utility trade-offs in mechanism design. Maximal leakage, hypothesis testing privacy, and other relaxations further broaden the analytic toolkit for measuring privacy risk.
Privacy-Utility Trade-off. Balancing the trade-off between privacy and utility is a central challenge in the design of privacy-preserving mechanisms. This challenge is frequently formulated as an optimization problem [30, 40, 32, 19, 22, 18, 12, 1, 20]. For example, Ghosh et al. [19] demonstrated that the geometric mechanism is universally optimal for differential privacy under certain loss-minimizing criteria in Bayesian settings, while Lebanon et al. [30] and Alghamdi et al. [1] studied utility-constrained optimization. Gupte et al. [22] modeled the privacy-utility trade-off as a zero-sum game between privacy mechanism designers and adversaries, illustrating the interplay between optimal privacy protection and worst-case loss minimization.
@@
2 Preliminaries
2.1 PAC Privacy
Privacy Threat Model. We consider the following general privacy problem. A sensitive input $X$ (e.g., a dataset, membership status) is drawn from a distribution $\mathcal{D}$, which may be unknown or inaccessible. There is a data processing (possibly randomized) mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}\subset\mathbb{R}^{d}$, where $\mathcal{Y}$ is measurable. An adversary observes the output $Y=\mathcal{M}(X)$ and attempts to estimate the original input $X$ with an estimate $\widetilde{X}$. The adversary has complete knowledge of both the data distribution $\mathcal{D}$ and the mechanism $\mathcal{M}$, representing the worst-case scenario. The central privacy concern is determining whether the adversary can accurately estimate the true input, meeting some predefined success criterion captured by an indicator function $\rho$.
The PAC privacy framework [46] addresses this threat model and is formally defined as follows.
@@ 2.1 PAC Privacy
For a data processing mechanism $\mathcal{M}$, given some data distribution $\mathcal{D}$, and a measure function $\rho(\cdot,\cdot)$, we say $\mathcal{M}$ satisfies $(\delta,\rho,\mathcal{D})$-PAC Privacy if the following experiment is impossible:
A user generates data $X$ from distribution $\mathcal{D}$ and sends $\mathcal{M}(X)$ to an adversary. The adversary who knows $\mathcal{D}$ and $\mathcal{M}$ is asked to return an estimation $\widetilde{X}\in\mathcal{X}$ on $X$ such that with probability at least $1-\delta$, $\rho(\widetilde{X},X)=1$.
@@
A mechanism $\mathcal{M}$ is termed $(\Delta_{f}^{\delta},\rho,\mathcal{D})$ PAC-advantage private if it is $(\delta,\rho,\mathcal{D})$ PAC private and
$\Delta_{f}^{\delta}\equiv\mathcal{D}_{f}(\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}})=\delta^{\rho}_{o}f(\frac{\delta}{\delta^{\rho}_{o}})+(1-\delta^{\rho}_{o})f(\frac{1-\delta}{1-\delta^{\rho}_{o}}).$
@@
Here, PAC Advantage Privacy is defined on top of PAC Privacy and quantifies the amount of privacy loss incurred from releasing $\mathcal{M}(X)$, captured by the additional posterior advantage $\Delta_{f}^{\delta}$.
@@ 2.2 Automatic PAC Privatization Algorithms
PAC Privacy enables automatic privatization, which supports simulation-based implementation for arbitrary black-box mechanisms, without requiring the worst-case adversarial analysis, such as sensitivity computation. In this section, we present the main theorems and algorithms underlying automatic PAC privatization as introduced in [46] (hereafter "Auto-PAC") and the efficiency-improved version proposed in [43] (hereafter "Efficient-PAC"; algorithm details in Appendix A). We start by defining the mutual information.
@@
$\mathtt{MI}(x;w)\equiv\mathcal{D}_{KL}(\mathsf{P}_{x,w}\|\mathsf{P}_{x}\otimes\mathsf{P}_{w}),$
@@
When the $f$-divergence in $\Delta_{f}^{\delta}$ is instantiated as the KL divergence (denoted as $\Delta_{\mathrm{KL}}^{\delta}$), Theorem 1 of [46] shows
$\Delta_{\mathrm{KL}}^{\delta}\leq\mathtt{MI}(X;\mathcal{M}(X)),$ (1)
Thus, one can control the posterior advantage $\Delta_{\mathrm{KL}}^{\delta}$ by bounding the mutual information between private data and the released output.
Next, we introduce the Auto-PAC. Consider a deterministic data processing mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where the output norm is uniformly bounded: $\|\mathcal{M}(X)\|_{2}\leq r$ for all $X$. To ensure PAC Privacy, the mechanism is perturbed by Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$, where $\Sigma_{B}$ is the covariance. For any deterministic mechanism $\mathcal{M}$ and any Gaussian noise $B$, define the Gaussian surrogate bound
@@
where $\Sigma_{\mathcal{M}(X)}$ is the covariance of $\mathcal{M}(X)$.
@@
For an arbitrary deterministic mechanism $\mathcal{M}$ and Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$, the mutual information satisfies
@@
Algorithm 1 $(1-\gamma)$-Confidence Auto-PAC [46]
0: deterministic mechanism $M$, dataset $\mathcal{D}$, sample size $m$, security parameter $c$, mutual information quantities $\beta^{\prime}$ and $v$.